OpenStudy (anonymous):

Match the prefix to the proper number associated with it. Match Term Definition deci A) .1 centi B) .01 kilo C) .001 milli D) 1000

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

Kilo=1000 or 1k Milli=0.001 Deci=10/100 or 0.1 Centi=.01

OpenStudy (dtan5457):

Kilo = 10^3 Hecto = 10^2 Deka = 10^1 Deci = 10^-1 Centi = 10^-2 Milli = 10^-3 Micro = 10^-6

OpenStudy (anonymous):

thanks :)